OpenStudy (anonymous):

3-2(x-1)=2+4x find x

OpenStudy (anonymous):

distribute 2 into (x-1) first

OpenStudy (anonymous):

okay so im done with all the steps except the last part where u divide im having trouble there

OpenStudy (anonymous):

there is no dividing needed, just adding and subtracting

OpenStudy (anonymous):

3-2x+2=2+4x distribute 5-2x=2+4x ad like terms namely the 2 and 3 in LHS 3-2x=4x sub 2 3=6x add 2x 3/6=x div by 6 1/2=x

OpenStudy (anonymous):

last step you dividide @monkeywall101

OpenStudy (anonymous):

thanks thats waht i gott i waz so confused when @monkeywall101 said no dividing :)

OpenStudy (anonymous):

sorry i was trying to do it in my head

OpenStudy (anonymous):

lol :) me 2 but it always better to write it @monkeywall101

OpenStudy (anonymous):

yup i got that thanks :)

OpenStudy (anonymous):

yea thats true but thanks t all who answered the question love u all :)

OpenStudy (anonymous):

aww :P

OpenStudy (anonymous):

haha ur welcome :)